# Tensor Product Attention Is All You Need

### v3

@@ A.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{q}$):
$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{k}$):
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (A.1)
A.2 Multi-Head Case
For multi-head attention with $h$ heads, one repeats the factorization across all heads. The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)}$ vectors are shared across heads.
A.3 Complexity Analysis
@@
    Naïve: Materialize $\mathbf{Q}$ and $\mathbf{K}$ from factors, then perform the usual batched GEMM.
@@
    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from the rank-$(R_{q},R_{k})$ factors without explicitly forming $\mathbf{Q},\mathbf{K}$.
@@
  • Constructing factors: $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times R_{q}(h+d_{h})+R_{k}(h+d_{h})+R_{v}(h+d_{h})\bigr{)}$.
  • Materializing $\mathbf{Q},\mathbf{K}$: $\mathcal{O}\bigl{(}B\,T\,(R_{q}\,h\,d_{h}+R_{k}\,h\,d_{h})\bigr{)}$.
  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$: $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$.
Typically $R_{q},R_{k},R_{v}\ll h$, so the overhead of constructing factors is small relative to $\mathcal{O}(T^{2}\,d_{\text{model}})$. Meanwhile, we still gain KV caching benefits.
TPA: Specialized Implementation. If we bypass explicitly forming $\mathbf{Q},\mathbf{K}$, each dot product $\mathbf{Q}_{t}\cdot\mathbf{K}_{\tau}$ is a double sum over rank indices. Below we detail its complexity.
A.4 Complexity Analysis for the Specialized Implementation
Single-Head Complexity. A single attention head of dimension $d_{h}$. For each query:
$\mathbf{Q}^{(i)}_{t}~{}=~{}\frac{1}{R_{q}}\sum_{r=1}^{R_{q}}a_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
and for each key:
$\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{k}}\sum_{s=1}^{R_{k}}a_{k,i}^{(s)}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
Their dot product:
$\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}~{}=~{}\frac{1}{R_{q}R_{k}}\sum_{r=1}^{R_{q}}\sum_{s=1}^{R_{k}}\Bigl{[}a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})\Bigr{]}~{}\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$
For each pair $(r,s)$, we pay:
@@ A.4 Complexity Analysis for the Specialized Implementation
    $\mathcal{O}(1)$ for multiplying two scalars,
@@
    $\mathcal{O}(d_{h})$ for the dot product $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})\cdot\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$.
Since $(r,s)$ runs over $R_{q}\times R_{k}$, each token-pair $(t,\tau)$ costs roughly
$\mathcal{O}\Bigl{(}R_{q}\,R_{k}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}~{}\approx~{}\mathcal{O}(R_{q}\,R_{k}\,d_{h}).$
For $T$ queries and $T$ keys, that is $\mathcal{O}(T^{2}\,R_{q}\,R_{k}\,d_{h})$ for a single head.
Multi-Head and Batches (Reusing $\mathbf{b}$-Dot Products). When extending to $h$ heads, each head $i$ has its own scalar factors $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$, but the $\mathbf{b}$-vectors $\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t})$ and $\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})$ can still be shared across all heads (assuming the same rank-$R$ factors for every head). Hence, one can split the total cost into two stages:
@@
    $\mathbf{b}$-Dot-Product Stage:
    For each token pair $(t,\tau)$ and each rank pair $(r,s)$, compute the dot product
    $\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),~{}\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}~{}\in~{}\mathbb{R}.$
    Since each dot product is $\mathcal{O}(d_{h})$ and there are $R_{q}R_{k}$ rank pairs as well as $T^{2}$ token pairs, this stage costs:
    $\mathcal{O}\bigl{(}T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}.$
    Crucially, these $\mathbf{b}$-dot products need only be computed once and can be cached for reuse by all heads.
@@
    Per-Head Scalar Multiplications:
    After the $\mathbf{b}$-dot products are precomputed (and cached), each head $i$ only needs to multiply each stored dot product by the corresponding scalars $\mathbf{a}_{q,i}^{(r)}(\mathbf{x}_{t})\,\mathbf{a}_{k,i}^{(s)}(\mathbf{x}_{\tau})$. Since this scalar multiplication is $\mathcal{O}(1)$ per pair, and there are $T^{2}$ token pairs and $R_{q}R_{k}$ rank pairs for each of the $h$ heads, this step costs:
    $\mathcal{O}\bigl{(}h\,T^{2}\,R_{q}R_{k}\bigr{)}.$
@@
Putting these together, for batch size $B$, the total cost is
$\mathcal{O}\bigl{(}B\,T^{2}\,R_{q}R_{k}\,d_{h}\bigr{)}~{}+~{}\mathcal{O}\bigl{(}B\,T^{2}\,h\,R_{q}R_{k}\bigr{)}~{}=~{}\mathcal{O}\Bigl{(}B\,T^{2}\,R_{q}R_{k}\bigl{(}d_{h}+h\bigr{)}\Bigr{)}.$
By contrast, the standard multi-head attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,h\,d_{h}\bigr{)}$. Hence, for the specialized TPA approach to reduce flops,
$R_{q}R_{k}\,(d_{h}+h)~{}\leq~{}h\,d_{h}.$
Thus a practical guideline is to ensure $\,R_{q}R_{k}<h\,\tfrac{d_{h}}{d_{h}+h}\,$. When that holds, bypassing explicit materialization of $\mathbf{Q}$ and $\mathbf{K}$ can be beneficial.
A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
We have explored a two-step procedure for computing $\mathbf{Q}\,\mathbf{K}^{\top}$ directly from factorized queries and keys without materializing $\mathbf{Q}$ or $\mathbf{K}$. Here, we extend this idea to also avoid explicitly forming $\mathbf{V}$. That is, all three activations $\mathbf{Q},\mathbf{K},\mathbf{V}$ remain factorized throughout the attention pipeline. We present a single-head formulation below, and then discuss multi-head and batch extensions.
@@ A.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
Extending the Two-Step Approach to Avoid $\mathbf{V}$ Materialization. After we obtain $\mathbf{Q}\mathbf{K}^{\top}$, we apply $\alpha_{t,\tau}=\operatorname{softmax}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,(QK^{\top})_{t,\tau}\bigr{)}$. The final attention output at token $t$ (single head) is
$\operatorname{head}(t)~{}=~{}\sum_{\tau=1}^{T}\alpha_{t,\tau}~{}\mathbf{V}_{\tau}.$
Using the factorization $\mathbf{V}_{\tau}=\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$, we write:
$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
Rearrange sums:
@@
$\operatorname{head}(t)~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
We still do not explicitly form $\mathbf{V}_{\tau}$. Instead:
Stage 1: Calculating $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for all tokens. We simply observe that each output $\operatorname{head}(t)$ can be computed by summing vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$ weighted by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. The complexity for constructing $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ $\forall u,\tau$ is $\mathcal{O}(T\,R_{v}\,d_{h})$.
Stage 2: Weighted Summation by $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. For each token $t$, the final attention head output is
$\frac{1}{R_{v}}\sum_{\tau=1}^{T}\alpha_{t,\tau}\sum_{u=1}^{R_{v}}a_{v}^{(u)}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})~{}=~{}\frac{1}{R_{v}}\sum_{u=1}^{R_{v}}\Bigl{[}\sum_{\tau=1}^{T}\bigl{(}\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\Bigr{]}.$
We still never explicitly materialize $\mathbf{V}$. Instead, for each pair $(t,u)$, we must accumulate the sum of $T$ vectors $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})\in\mathbb{R}^{d_{h}}$, each scaled by the scalar $\alpha_{t,\tau}\,a_{v}^{(u)}(\mathbf{x}_{\tau})$. Because each vector is $d_{h}$-dimensional, each $(t,u)$ summation costs $\mathcal{O}(T\,d_{h})$. Summed over $t=1\ldots T$ and $u=1\ldots R_{v},$ the total work is $\mathcal{O}(T^{2}\,R_{v}\,d_{h})$ for the entire sequence.
In practice, one precomputes all $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau})$ for $\tau=1\ldots N$, so each accumulation can be implemented as a simple “scalar‐times‐vector add” in a tight loop. This cost is usually smaller than the $\mathbf{Q}\mathbf{K}^{\top}$ factorized cost if $R_{v}\ll h$.
A.6 Overall Complexity for Single-Head
Combining the four bullet‐point stages from above (ignoring smaller overheads like the softmax) yields:

### v4

@@ C.1 Single-Head Factorization Setup Without Materializing $\mathbf{Q}$ and $\mathbf{K}$
Consider a single head $i$. Each query vector $\mathbf{Q}^{(i)}_{t}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{Q}$ and scale $s_{Q}=1/R_{Q}$):
$\mathbf{Q}^{(i)}_{t}~{}=~{}s_{Q}\sum_{r=1}^{R_{Q}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),$
@@
and each key vector $\mathbf{K}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{K}$ and scale $s_{K}=1/R_{K}$):
$\mathbf{K}^{(i)}_{\tau}~{}=~{}s_{K}\sum_{s=1}^{R_{K}}a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau}).$
@@
$\bigl{[}\mathbf{Q}^{(i)}\,(\mathbf{K}^{(i)})^{\top}\bigr{]}_{t,\tau}~{}=~{}s_{Q}s_{K}\sum_{r=1}^{R_{Q}}\sum_{s=1}^{R_{K}}a^{(r)}_{q,i}(\mathbf{x}_{t})\,a^{(s)}_{k,i}(\mathbf{x}_{\tau})\,\bigl{\langle}\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\bigr{\rangle}.$ (C.1)
Similarly, the value vector $\mathbf{V}^{(i)}_{\tau}\in\mathbb{R}^{d_{h}}$ is factorized (with rank $R_{V}$ and scale $s_{V}=1/R_{V}$):
$\mathbf{V}^{(i)}_{\tau}~{}=~{}s_{V}\sum_{u=1}^{R_{V}}a^{(u)}_{v,i}(\mathbf{x}_{\tau})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau}).$
C.2 Multi-Head Case
For multi-head attention with $h$ heads, one repeats the factorization across all heads. The $\mathbf{b}_{q}^{(r)},\mathbf{b}_{k}^{(s)},\mathbf{b}_{v}^{(u)}$ vectors are shared across heads. The scaling factors $s_{Q},s_{K},s_{V}$ are applied per head as part of the definition.
C.3 Complexity Analysis
@@
    Naïve: Materialize $\mathbf{Q},\mathbf{K},\mathbf{V}$ from factors (including their scales), then perform the usual batched GEMM.
@@
    Specialized: Attempt to compute $\mathbf{Q}\,\mathbf{K}^{\top}$ and the final attention output directly from the rank-$(R_{Q},R_{K},R_{V})$ factors and their scales, without explicitly forming full $\mathbf{Q},\mathbf{K},\mathbf{V}$.
@@
  • Constructing factors:   $\mathcal{O}\bigl{(}B\,T\,d_{\text{model}}\,\times(R_{Q}(h+d_{h})+R_{K}(h+d_{h})+R_{V}(h+d_{h}))\bigr{)}$.
  • Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$ (including scaling):   $\mathcal{O}\bigl{(}B\,T\,(R_{Q}\,h\,d_{h}+R_{K}\,h\,d_{h}+R_{V}\,h\,d_{h})\bigr{)}$. The scaling is a per-element multiplication.
  • Dot-product $\mathbf{Q}\,(\mathbf{K})^{\top}$ and multiply by $\mathbf{V}$:   $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for scores, and $\mathcal{O}\bigl{(}B\,T^{2}\,d_{\text{model}}\bigr{)}$ for value aggregation.
Typically $R_{Q},R_{K},R_{V}\ll h,d_{h}$, so the overhead of constructing factors and materializing is influenced by these ranks. The key benefit of TPA is in reducing KV cache size and potentially computation if specialized kernels are used.
TPA: Specialized Implementation. If we bypass explicitly forming $\mathbf{Q},\mathbf{K},\mathbf{V}$, the computation involves the sums over ranks and the scaling factors. Below, we detail its complexity.
Algorithm 2 Specialized TPA Computation (Without Materializing $\mathbf{Q},\mathbf{K},\mathbf{V}$)
1:Factorized Query components: $\mathbf{A}_{Q}(T,H,R_{Q})$, $\mathbf{B}_{Q}(T,R_{Q},d_{h})$
2:Factorized Key components: $\mathbf{A}_{K}(T,H,R_{K})$, $\mathbf{B}_{K}(T,R_{K},d_{h})$
3:Factorized Value components: $\mathbf{A}_{V}(T,H,R_{V})$, $\mathbf{B}_{V}(T,R_{V},d_{v})$ (Note: $d_{v}$ is value head dim, often $d_{v}=d_{h}$)
4:Sequence length $T$, Number of heads $H$, Ranks $R_{Q},R_{K},R_{V}$, Head feature dimension $d_{h}$
5:Scaling factors: $s_{Q}=1/R_{Q},s_{K}=1/R_{K},s_{V}=1/R_{V}$ (or other pre-defined scales)
6:Output tensor $\mathbf{O}(T,H,d_{v})$
7:// Phase 1: Attention Score Computation
8$\triangleright$ Step 1.1: Compute shared $\mathbf{B}$-factor dot products $P$
9:$P\leftarrow\text{einsum}(\text{"tqrd, tkrd}\rightarrow\text{tqtkrqrk"},\mathbf{B}_{Q},\mathbf{B}_{K})$
10$\triangleright$ Indices: $t_{q}$=query token, $t_{k}$=key token, $r_{q}$=query rank, $r_{k}$=key rank, $d$=feature dim
11$\triangleright$ Step 1.2: Combine $P$ with $\mathbf{A}_{Q},\mathbf{A}_{K}$ and apply $s_{Q},s_{K}$ to form scaled logits $\mathcal{L}^{\prime}$
12:$\mathcal{L}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqhrq, tkhrk, tqtkrqrk}\rightarrow\text{tqtkh"},\mathbf{A}_{Q},\mathbf{A}_{K},P)$
13:$\mathcal{L}^{\prime}\leftarrow s_{Q}\cdot s_{K}\cdot\mathcal{L}_{\text{unscaled}}$
14$\triangleright$ Indices: $h$=head
15$\triangleright$ Step 1.3: Apply attention scaling ($1/\sqrt{d_{h}}$) and Softmax to get probabilities $\bm{\alpha}$
16:$\mathcal{L}^{\prime\prime}_{\text{scaled}}\leftarrow\mathcal{L}^{\prime}/\sqrt{d_{h}}$
17:$\bm{\alpha}\leftarrow\operatorname{Softmax}_{T_{k}}(\mathcal{L}^{\prime\prime}_{\text{scaled}})$triangleright Softmax over key tokens $T_{k}$ for each query $T_{q}$ and head $H$
18$\triangleright$ Output $\bm{\alpha}(t_{q},t_{k},h)$
19:
20:// Phase 2: Value Aggregation
21$\triangleright$ Step 2.1: Conceptually weight $\mathbf{A}_{V}$ by $\bm{\alpha}$ to form $W_{\mathbf{A}_{V}}$
22:$W_{\mathbf{A}_{V}}\leftarrow\text{einsum}(\text{"tqtkh, tkhrv}\rightarrow\text{tqtkhrv"},\bm{\alpha},\mathbf{A}_{V})$
23$\triangleright$ Indices: $r_{v}$=value rank
24$\triangleright$ Step 2.2: Aggregate $W_{\mathbf{A}_{V}}$ with $\mathbf{B}_{V}$ and apply $s_{V}$ to get final output $\mathbf{O}$
25:$\mathbf{O}_{\text{unscaled}}\leftarrow\text{einsum}(\text{"tqtkhrv, tkrvd}\rightarrow\text{tqhd"},W_{\mathbf{A}_{V}},\mathbf{B}_{V})$triangleright $d$ is value feature dim $d_{v}$
26:$\mathbf{O}\leftarrow s_{V}\cdot\mathbf{O}_{\text{unscaled}}$
27:return $\mathbf{O}$
C.4 Complexity Analysis for the Specialized Implementation
Complexity for a Single Query Token (Autoregressive Decoding). The dot product $\mathbf{Q}^{(i)}_{t}\cdot\mathbf{K}^{(i)}_{\tau}$ from Eq. (C.1) includes the $s_{Q}s_{K}$ scaling. For each pair $(r,s)$, we pay:
@@ C.4 Complexity Analysis for the Specialized Implementation
    $\mathcal{O}(1)$ for multiplying $a_{q,i}^{(r)}(\mathbf{x}_{t})\,a_{k,i}^{(s)}(\mathbf{x}_{\tau})$,
@@
    $\mathcal{O}(d_{h})$ for the dot product $\langle\mathbf{b}_{q}^{(r)}(\mathbf{x}_{t}),\mathbf{b}_{k}^{(s)}(\mathbf{x}_{\tau})\rangle$.
The $s_{Q}s_{K}$ scaling is a single multiplication after the sums. Since $(r,s)$ runs over $R_{Q}\times R_{K}$, each token-pair $(t,\tau)$ costs roughly $\mathcal{O}\Bigl{(}R_{Q}\,R_{K}\,\bigl{(}1+d_{h}\bigr{)}\Bigr{)}\approx\mathcal{O}(R_{Q}\,R_{K}\,d_{h})$. For a single query token $t$ and $M$ cached key tokens $\tau=1\dots M$, this is $\mathcal{O}(M\,R_{Q}\,R_{K}\,d_{h})$ for a single head to compute all unscaled dot products. The scaling by $s_{Q}s_{K}$ is $\mathcal{O}(M)$ for the score vector.
Multi-Head and Batches (Full Sequence Processing - reusing $\mathbf{b}$-Dot Products as in Algorithm 2). The $\mathbf{b}$-vectors can be shared across heads. Let $T$ be the sequence length.
@@
    $\mathbf{b}$-Dot-Product Stage ($P$):
    Compute $P(t_{q},\tau_{k},r_{q},r_{k})=\langle\mathbf{b}_{q}^{(r_{q})}(\mathbf{x}_{t_{q}}),\mathbf{b}_{k}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\rangle$. Cost: $\mathcal{O}\bigl{(}T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}$. This is shared across heads.
@@
    Per-Head Score Computation:
    For each head $h$, compute $\mathcal{L}_{t_{q},\tau_{k},h}=s_{Q}s_{K}\sum_{r_{q},r_{k}}a_{q,h}^{(r_{q})}(\mathbf{x}_{t_{q}})\,a_{k,h}^{(r_{k})}(\mathbf{x}_{\tau_{k}})\,P(t_{q},\tau_{k},r_{q},r_{k})$.
    Cost: $\mathcal{O}\bigl{(}H\,T^{2}\,R_{Q}R_{K}\bigr{)}$.
@@
Putting these together for batch size $B$, the total cost for scores is
$\mathcal{O}\bigl{(}B\,T^{2}\,R_{Q}R_{K}\,d_{h}\bigr{)}+\mathcal{O}\bigl{(}B\,H\,T^{2}\,R_{Q}R_{K}\bigr{)}=\mathcal{O}\Bigl{(}B\,T^{2}\,R_{Q}R_{K}\bigl{(}d_{h}+H\bigr{)}\Bigr{)}$. This is for computing the scaled logits before softmax. The standard attention dot-product step is $\mathcal{O}\bigl{(}B\,T^{2}\,H\,d_{h}\bigr{)}$. For the specialized TPA to reduce flops in score computation: $R_{Q}R_{K}\,(d_{h}+H)<H\,d_{h}$. This implies $R_{Q}R_{K}<H/(1+H/d_{h})$. If $H\approx d_{h}$, then $R_{Q}R_{K}\approx H/2$.
C.5 Toward Faster Computation Without Materializing $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$
@@
We extend the idea to also avoid explicitly forming $\mathbf{V}$. After obtaining scaled logits $\mathcal{L}^{\prime}_{t_{q},\tau_{k}}=s_{Q}s_{K}(\mathbf{Q}_{t_{q}}\mathbf{K}_{\tau_{k}}^{\top})$ (per head), we apply $\alpha_{t_{q},\tau_{k}}=\operatorname{softmax}_{\tau_{k}}\bigl{(}\tfrac{1}{\sqrt{d_{h}}}\,\mathcal{L}^{\prime}_{t_{q},\tau_{k}}\bigr{)}$. The final attention output at token $t_{q}$ (single head) using $\mathbf{V}_{\tau_{k}}=s_{V}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ is:
$\operatorname{head}(t_{q})~{}=~{}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}~{}\mathbf{V}_{\tau_{k}}~{}=~{}s_{V}\sum_{\tau_{k}=1}^{T}\alpha_{t_{q},\tau_{k}}\sum_{u=1}^{R_{V}}a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}}).$
Rearranging sums:
@@
$\operatorname{head}(t_{q})~{}=~{}s_{V}\sum_{u=1}^{R_{V}}\left[\sum_{\tau_{k}=1}^{T}\bigl{(}\alpha_{t_{q},\tau_{k}}\,a_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\bigr{)}\,\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})\right].$
The computation of $\mathbf{b}_{v}^{(u)}(\mathbf{x}_{\tau_{k}})$ from $\mathbf{x}_{\tau_{k}}$ is $\mathcal{O}(T\,R_{V}\,d_{\text{model}}d_{v})$ if contextual, or loaded if fixed (where $d_{v}$ is the dimension of $\mathbf{b}_{v}$). Assuming $\mathbf{b}_{v}^{(u)}$ are available, the weighted summation for each $(t_{q},u)$ pair (inner sum over $\tau_{k}$) costs $\mathcal{O}(T\,d_{v})$. Summed over $t_{q}=1\ldots T$ and $u=1\ldots R_{V}$, this stage is $\mathcal{O}(T^{2}\,R_{V}\,d_{v})$. The final scaling by $s_{V}$ is $\mathcal{O}(Td_{v})$.
C.6 Overall Complexity for Single-Head (Specialized)
The dominant costs for a single head, assuming factors $\mathbf{A},\mathbf{B}$ are already computed from $\mathbf{X}$: